### v2

# Topological Weyl Altermagnetism in CrSb

Abstract
Altermagnets constitute a novel, third fundamental class of collinear magnetic ordered materials, alongside with ferro- and antiferromagnets [1, 2]. They share with conventional antiferromagnets the feature of a vanishing net magnetization. At the same time they show a spin-splitting of electronic bands, just as in ferromagnets, caused by the atomic exchange interaction[3, 4, 5]. On the other hand, topology has recently revolutionized our understanding of condensed matter physics, introducing new phases of matter classified by intrinsic topological order[6, 7, 8]. Here we connect the worlds of altermagnetism and topology, showing that the electronic structure of the altermagnet CrSb is topological and hosts a novel Weyl semimetallic state. Using high-resolution and spin angle-resolved photoemission spectroscopy, we observe a large momentum-dependent spin-splitting in CrSb, reaching up to 1 eV, that induces altermagnetic Weyl nodes with an associated magnetic quantum number. At the surface we observe their spin-polarized topological Fermi-arcs. This establishes that in altermagnets the large energy scale intrinsic to the spin-splitting – orders of magnitude larger than the relativistic spin-orbit coupling – creates its own realm of robust electronic topology.
The recently established new class of altermagnetic materials is characterized by the presence of particular alternating opposite-spin sublattices. They defy conventional magnetic classifications by breaking time-reversal symmetry while simultaneously maintaining a zero net magnetization enforced by spin-lattice symmetry[1, 2]. Precisely this intrinsic trait allows altermagnets to combine advantageous features of both ferromagnets and antiferromagnets[9, 10], including spin-transport properties and compatibility with a diverse range of materials[11, 12, 13, 14, 15, 16, 17, 18], from superconductors[19, 20, 22, 23, 21] to insulators[24, 25, 26]. As the vanishing net magnetization in altermagnets avoids effects of long-range magnetic stray fields typical for ferromagnets, this combination opens up promising avenues for applications in magneto-optics[27], spintronics[28, 14], and beyond.
Interestingly, symmetry analysis reveals that the electronic structure of altermagnets may in principle provide fertile ground for non-trivial electronic topology, due to the unique interplay between time-reversal and crystalline symmetries[30, 29, 31, 32, 33]. Despite extensive experimental and theoretical scrutiny of altermagnets[25, 35, 24, 26, 36, 37, 38, 34, 39, 40, 41] and topological semimetals[42, 6, 43, 44, 45, 46, 47, 48, 6, 43] separately, the simultaneous experimental manifestation of their characteristic phenomena has remained elusive so far. Here we establish CrSb as a topological Weyl semimetal with two distinct types of spin-carrying nodal structures. The giant spin splitting that we find from high resolution angle-resolved photoemission spectroscopy (ARPES) and spin-ARPES measurements aligns well with density functional theory (DFT) calculations on CrSb, as do the bulk Weyl points (WPs) with linear dispersion and ensuing altermagnetic surface Fermi arcs (SFAs) that, as we will show, connect same-spin WPs at the surface.
For both its altermagnetic and topological features, the symmetries of CrSb are crucial. Figure 1a shows the crystal and magnetic structure of CrSb, which forms a hexagonal structure with the space group $P6_{3}/mmc$ (No. 194)[49]. The magnetic space group is $P6^{\prime}_{3}/m^{\prime}m^{\prime}c$ (No. 194.268) and the spin space group is $P^{-1}{6_{3}/}^{-1}{m}^{1}{m}^{-1}c^{\infty m}1$ with a six-fold screw rotation symmetry connecting Cr-atoms of opposite magnetic moment in real space. The three-dimensional (3D) Fermi surface of CrSb from DFT calculations in Fig. 1c illustrates that sixfold rotation connecting opposite-spin sublattices in real space (Fig. 1a) also connects opposite-spin electronic states in momentum space (Fig. 1c), as is mandatory for altermagnets[1, 2]. The calculated Fermi surface also illustrates the four mirror planes in the Brillouin zone (BZ) on which the altermagnetic bands are spin-degenerate in absence of spin-orbit coupling (SOC), as shown in Fig. 1b.
The altermagnetic spin-splitting allows for topological electronic structures far beyond antiferromagnets with completely spin-degenerate bands[30, 29, 31, 32, 33]. We identify two mechanisms for the generation of topological nodes in the electronic structure of CrSb. The first relies on the observation that spin-splitting of fully spin-polarized altermagnetic bands automatically leads to nodal lines when opposite-spin bands cross. The degenerate states that form the nodal lines have opposite spins. Such crossings in momentum space can be symmetry enforced, for instance at a mirror plane, or be accidental. The effect of SOC in CrSb is to lift this spin-degeneracy for most, but not all momenta on nodal lines. At the remaining Weyl nodes also the spin degeneracy remains, so that the points correspond to doublets with vanishing total spin-projection $S^{z}=0$ as illustrated in Fig. 1d. The altermagnetic symmetries and resulting spin-degeneracies thus provide a very natural setting for such opposite-spin WPs governed by the SOC energy scale. Indeed from DFT we identify 10 groups of opposite-spin WPs within 1 eV from the Fermi energy, see supplementary material for details.
Interestingly, there is also a route that generates purely altermagnetic WPs, reflecting the bulk altermagnetic symmetries. They give rise to same-spin WPs with a total spin-projection of $S^{z}=\pm 1$ (in units of $\hbar$) and a spin-splitting (energy-splitting at the nodal momentum) governed by the exchange interaction energy scale[1, 2], as illustrated in Fig. 1e. These altermagnetic Weyl nodes thus carry apart from a topological charge also a magnetic quantum number $S^{z}$, and with it, a finite magnetic moment. It is precisely the altermagnetic crystal symmetry that relates same-spin WPs with different $S^{z}$. The weak SOC that formally breaks the altermagnetic symmetry in CrSb causes the expectation value $|\langle S^{z}\rangle|$ to slightly deviate from unity (see supplemental material). These WPs, and the topological surface states they imply, render the topological features of Weyl altermagnets fundamentally distinct from ferro- or antiferromagnetic Weyl semimetals.
For CrSb, where lattice inversion symmetry connects atoms with the same magnetization, the supplementary material provides details of the topological same-spin nodal line without SOC and three groups of same-spin WPs that we identified. Fig. 1f- 1g shows the distribution of same-spin WPs in momentum space, with their spin-projection quantum number. Fermi-arcs connect the surface projections of these WPs of opposite chirality. As a consequence also their topological surface states are fully spin-polarized in the altermagnetic limit, with additional SOC affecting the details of their dispersion and spectral weight. The electronic structures of both Sb (Fig. 1h) and Cr (Fig. 1i) terminated surfaces show that the surface states are very extended in momentum space and carry substantial spectral weight. Detailed representations of purely surface related states and the full 3D bulk band structures are provided in the supplementary material.
To experimentally access the spin-splitting and topological electronic structure of CrSb we use ARPES, which is to a certain extent complicated by CrSb having a 3D electronic structure. In order to establish the correspondence between photon energy and out-of-plane momentum kz, we performed broad-range (40 to 120 eV) photon energy dependent ARPES measurements along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. Fig. 2i shows the photon energy dependent ARPES spectral intensity map (kx-kz Fermi surface) at the Fermi level along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. From the periodic structure along the kz direction, the correspondence between the high symmetry points of the BZ along the kz direction and the photon energy is determined as shown in Fig. 2i. To distinguish between bulk and surface states, we measured Fermi surfaces on a relatively uneven sample surface for different kz, corresponding to photon energies from 97 eV to 122 eV (Fig. 2k). Since the uneven surface suppresses the contribution of surface states[47, 50], these Fermi surfaces mainly derive from the bulk. All their features are captured by the calculated bulk Fermi surface with corresponding kz (Fig. 2c- 2h). The Fermi surfaces away from the kz = 0 $\pi/c$ and 1 $\pi/c$ planes (Fig. 2d- 2f) have the form of a hexagram consisting of two intersecting equilateral triangular structures. Due to the altermagnetic order in CrSb, the two Fermi surfaces corresponding to the two different triangles, which are very clearly resolved experimentally, have opposite-spin polarizations (Fig. 2d- 2f). Interestingly, the Sb and Cr terminated surfaces (Fig. 2l and  2m) show additional features apart from the bulk Fermi surface, which are marked by orange arrows. As we will detail later, comparison to the calculated surface spectral density (Fig. 2a- 2b) establishes that their root cause are the topological SFAs that in turn originate from the WPs (see Fig. S9 in supplementary material for a further comparison of constant energy contours).
First we establish that there is a very large momentum dependent spin-splitting of the bulk bands in CrSb[1, 5, 36, 39, 40, 41]. Fig. 3a- 3c show the photon energy dependent dispersions along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction measured with photon energies from 97 eV to 122 eV with the interval of 5 eV for an uneven sample surface (Fig. 3a) and flat Sb (Fig. 3b) and Cr (Fig. 3c) terminated surfaces. The corresponding kz dependent calculated spin-polarized band structures are shown alongside it (Fig. 3d- 3i). These bands are spin degenerate on the kz = 0 $\pi/c$ and 1 $\pi/c$ planes. Due to the altermagnetism in CrSb the bands between these two kz planes (Fig. 3e- 3h) exhibit spin splitting. The calculated bands and their spin splittings are in very good agreement with the ARPES measurements [Fig. 3a- 3c, (ii-v)]. We determine the maximum energy band splitting to be up to 1 eV (see Fig. S10 in the supplementary material). To further verify that the split bands are spin polarized, we performed spin resolved ARPES measurements with a photon energy of 102 eV, as shown in Fig. 3j- 3l. Fig. 3j and  3k are the spin-polarized energy distribution curves (EDCs) along the left and right cut lines in Fig. 3a(ii). The spin-polarized intensity difference between spin-up and spin-down of left and right bands is shown in Fig. 3l. Due to the existence of a large non-spin polarized background, originating from a Shirley background, the measured spin polarization observed in the EDCs is limited. Despite this, a difference in energy-dependent spin polarization intensity between spin-up and spin-down of left and right bands (Fig. 3l) can still be observed. This indicates that the left and right energy bands [Fig. 3a(ii)] have different spin projections along the $z$ direction, in agreement with theory.
Having identified and quantified the altermagnetic spin-splitting in CrSb, we focus on the resulting topological properties. Both ARPES measurements and DFT calculations show that CrSb is metallic with a complex Fermi surface. In an energy range around from $-1$ eV to 1 eV around the Fermi level, we identify 13 groups of WPs in the calculated bands, as listed in Table. 1 of the supplementary material. All WPs in CrSb arise from time-reversal symmetry breaking and their relative positions are given by the symmetries in the magnetic space group $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268) reflecting their unique altermagnetic nature. Many WPs are along or near the $\rm\overline{\Gamma}-\overline{M}$ direction (see Fig. S2 in supplementary material) – this is the high-symmetry line whose little co-group allows non-zero chirality. The pair of WPs in the $\rm\overline{\Gamma}-\overline{M}$ direction below the Fermi level (marked by round dots in Fig. 4a- 4b and  4h- 4i) are same-spin WPs. Long SFAs connect neighboring WPs of opposite chirality $\chi=\pm 1$ and identical spin-projection across the BZ. This topological surface state happens to appear in a region without bulk states so that it may be clearly discerned in ARPES. Even if the presence of these SFAs is dictated by topology, the shape of the surface spectra for two different terminations differ due to the lack of inversion and mirror symmetry on the surface, as shown in Fig. 4. Due to the large distance between two WPs projected on the 001 surface, the arc is even detectable along $\rm\overline{\Gamma}-\overline{K}$ as shown in Fig. 4o- 4p.
To track these WPs and SFAs experimentally, we conducted photon energy dependent ARPES measurements along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}$ on Sb (Fig. 4c- 4g) and Cr (Fig. 4j- 4n) terminated surfaces. The band features measured on Sb (marked by orange arrows in Fig. 4c- 4g) and Cr (marked by orange arrows in Fig. 4j- 4n) surfaces exhibit negligible photon energy dependence [momentum distribution curve (MDC) analysis of band features measured on Cr terminated surface see Fig. S11 in supplementary material], evidencing that it corresponds to a surface state. Comparison with the corresponding calculated surface states, shows that their band features (marked by orange arrows in Fig. 4c- 4g and Fig. 4j- 4n) agree well with the SFAs (marked by orange arrows in Fig. 4b and  4i), indicating that the SFAs are intrinsic and robust. As pointed out earlier, the SFAs are also observed in the 2D momentum maps at the Fermi energy (Fig. 2l and  2m). Furthermore, the DFT calculations show the kz location of WPs at an energy of -0.357 eV (round dots in Fig. 4a- 4b and  4h- 4i) is $\pm$0.27 $\pi/c$. This corresponds to a photon energy of about 82 eV. We show the WP in the band dispersion panel measured with 82 eV (Fig. 4d) and indeed it is located right at the intersection of the bands. In addition, we also measured the band dispersion along $\rm\overline{\Gamma}-\overline{K}$ with photon energy of 85 eV on Sb (Fig. 4q) and Cr (Fig. 4t) surfaces (for more detailed photon energy dependent measurements along this direction see Fig. S12 in supplementary material). Also along these high-symmetry lines, SFAs (denoted by orange arrows in Fig. 4q and  4t) are observed for both surface terminations, reflecting the large portion of the surface BZ traversed by the SFA, consistent with the calculations.
Thus we have not only quantified the altermagnetic spin-spliting of bands in CrSb with high-resolution and spin-resolved ARPES measurements, but also established it as a topological Weyl semi-metal, with concomitant altermagnetic topological surface states. Whereas in general Weyl nodes carry only a topological charge, corresponding to their chirality $\chi$, in altermagnets they also carry spin-projection $S^{z}$. Besides opposite-spin nodes with zero spin-projection, altermagnetic CrSb also has same-spin nodes with spin-projection $\pm\hbar$ and chirality-spin locking. For each set of $n$ symmetry-related altermagnetic Weyl nodes one can define a locking invariant $\zeta=\sum_{i=1,n}\chi_{i}S^{z}_{i}$. For CrSb $\zeta=0$ by symmetry, but interestingly other magnetic symmetry groups allow sets of altermagnetic nodes with non-zero integer locking invariants. The distribution of same-spin nodes in 3D momentum space reflects all altermagnetic symmetries as do their spin-polarized topological surface states, the altermagnetic Fermi arcs. Spin-orbit coupling weakly breaks altermagnetic symmetries and consequently the quantization of the spin-projection. As these properties derive from the interplay of symmetry and topology, they are generic for Weyl altermagnets. These findings may well imply that the unique bulk altermagnetic spin-transport properties are promoted by topology to also become properties of the Fermi arcs at the surface, which can render same-spin Weyl altermagnets interesting interface materials for spintronics. These insights not only spotlight the distinctive altermagnetic attributes of CrSb but also its potential to induce novel physics and applications in the area of topological materials.
Methods
Sample
The CrSb single crystals were grown by the chemical vapor transport (CVT) method. A stoichiometric ratio of chromium and antimony powders, together with iodine of 2.5 mg/ml as the transport agent, were mixed and sealed in an evacuated quartz ampoule. The ampoule was slowly heated and finally exposed to a temperature gradient of 925°C to 900°C where the CVT preceded for one week, then naturally cooled down to room temperature. CrSb crystals in size of  5 mm with regular shapes and shiny surfaces were obtained.
ARPES Measurements High-resolution ARPES measurements were performed at the Bloch beamline of MAX IV and at the I05 beamline of the Diamond synchrotron light source. The total energy resolution (analyzer and beamline) was set at 15$\sim$20 meV for the measurements. The angular resolution of the analyser was $\sim$0.1 degree. The beamline spot size on the sample was about 10 $\mu$m×$\times$12 $\mu$m at the Bloch beamline of MAX IV and about 70 $\mu$m×$\times$70 $\mu$m at the I05 beamline of the Diamond synchrotron. The samples were cleaved in situ and measured at about 18 K at the Bloch beamline of MAX IV and about 10 K at the I05 beamline of the Diamond synchrotron in ultrahigh vacuum with a base pressure better than 1.0$\times$10-10 mbar. The spin-resolved ARPES (SARPES) measurements were performed at the Bloch B-branch beamline of MAX IV with photon energy of 102 eV. The samples were cleaved in situ and measured at about 77 K at the Bloch B-batch beamline of MAX IV. The total energy resolution (analyzer and beamline) was set at $\sim$50 meV for the measurements. The angular resolution of the analyser was $\sim$1.3 degree.
DFT calculations Full relativistic spin-polarized electronic structure and Fermi surface calculations are done using the full potential local orbital code FPLO [51, 52], on a $k$-mesh of $20\times 20\times 13$ points. We use the generalized gradient approximation to the exchange and correlation potential by Perdew, Burke, and Ernzerhof (PBE). The surface state and WPs calculations are performed based on the DFT calculation from VASP [53] employing the projector augmented wave method [54]. The Brillouin zone is sampled with a $10\times 10\times 8$, with Gamma-centered $k$-point. The energy cutoff of the plane wave basis is set to 550 eV. The Hubbard term was introduced and set to be 0.8 eV in the $d$ orbitals of Cr-atom in the DFT framework (DFT+U) in order to account for the electron-electron correlation. The Wannier based Hamiltonian is symmetrized based on the maximally localized Wannier functions generated by the WANNIER90 interface [55]. The projectors are $d$ orbitals of Cr and $p$ orbitals of Sb atoms with the well-fitted region from -2 to 2 eV. To locate the WPs and calculate the chirality, Wanniertools is implemented [56].
Data Availability
The authors declare that all data supporting the findings of this study are available within the paper and its Supplementary Information files.
Acknowledgements
The work presented here was financially supported by the Swedish Research council (2019-00701) and the Knut and Alice Wallenberg foundation (2018.0104). We thank Ulrike Nitzsche for technical assistance. We acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), through SFB 1143 (Project ID 247310070), project A05, Project No. 465000489, and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter, ct.qmat (EXC 2147, Project ID 390858490). M.G.V. thanks support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) GA3314/1-1 -FOR 5249 (QUAST) and to the Spanish Ministerio de Ciencia e Innovacion grant PID2022-142008NB-I0. M.L.H. thanks the support from the Alexander von Humboldt Foundation and the useful discussion and help of Iñigo Robredo. Z.L.L thanks the support from the Youth Innovation Promotion Association of Chinese Academy of Sciences (No. 2021008). H.L. was supported by the Swedish Research Council (VR, grant 2018-00313), the Wallenberg Academy Fellows program of the Knut and Alice Wallenberg Foundation (2018.0460) and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. We acknowledge MAX IV Laboratory for time on Beamline BLOCH under Proposal 20230262 and 20231119. Research conducted at MAX IV, a Swedish national user facility, is supported by the Swedish Research council under contract 2018-07152, the Swedish Governmental Agency for Innovation Systems under contract 2018-04969, and Formas under contract 2019-02496.
Author Contributions
C.L. and H.L. proposed the project. C.L., M.L.H., H.L., O.T. and J.V.D.B. conceived the project. C.L. carried out the ARPES experiments with the assistance from Y.W. and W.Y.C.. M.L.H. and O.J. contributed to the band structure calculations. Z.L.L. contributed to CrSb crystal growth. C.L. contributed to software development for data analysis and analyzed the data. C.L., M.L.H., H.L., O.T. and J.V.D.B. wrote the paper. B.T., M.L., C.P. and T.K. provided the beamline support. C.L, M.L.H., Y.W., H.L., C.F., M.V., O.J., O.T. and J.V.D.B. contributed to the scientific discussions. All authors participated in and commented on the paper.
Competing Interests
The authors declare no competing interests.
Figure 1: Crystal structure, WPs and calculated electronic structure of CrSb. (a) The crystal structure of CrSb with the space group $P6_{3}/mmc$ (no. 194). (b) The 3D BZ of the original unit cell of CrSb, and the corresponding two-dimensional BZ projected on the (001) plane (green lines). The red, green, blue and orange planes are the four mirror planes in the BZ on which the altermagnetic bands are spin-degenerate in absence of SOC. (c) Calculated bulk Fermi surface in the first BZ. The colormap from blue to orange of Fermi surface sheet represents the expectation value of spin along the $z$ direction. (d-e) Illustration of the two types of WPs: opposite-spin (d) and same-spin (e). Top (f) and side (g) views of the distribution of same-spin WPs in the 3D BZ. Red dots represent nodes with chirality $\chi=+1$, green dots $\chi=-1$; the spin polarizations are indicated by cyan and pink. The blue line represents spin-up and spin-down degeneracy. (h-i) The calculated surface state of (001) on Sb (h) and Cr (i) terminations along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}-\overline{K}_{2}-\overline{\Gamma}-\overline{K}_{1}-\overline{M}_{1}$. The WPs are marked by the red ($\chi=1$) and green ($\chi=-1$) dots. Star, square and round dots represent the opposite-spin WPs on the high symmetry line/plane (WP1), opposite-spin WPs on general momenta (WP2) and same-spin WPs (WP3), respectively.
Figure 2: Fermi surface of CrSb. (a-b) projected Fermi surfaces for Sb (a) and Cr (b) surface terminations with WPs overlapped. Red dots represent nodes with chirality $\chi=+1$, green dots $\chi=-1$; (c-h) Calculated bulk Fermi surfaces at kz = 1 $\pi/c$ (c), 0.8 $\pi/c$ (d), 0.6 $\pi/c$ (e), 0.4 $\pi/c$ (f), 0.2 $\pi/c$ (g) and 0 $\pi/c$ (h) planes, integrated over a ±$\pm 0.1$ $\pi/c$ interval. (i) Photon energy dependent ARPES spectral intensity map at Fermi level along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$. (j) corresponding calculated kz-kx Fermi surface. (k) Photon energy dependent Fermi surfaces measured with photon energies of 97 eV (i), 102 eV (ii), 107 eV (iii), 112 eV (iv), 117 eV (v), 122 eV (vi) on an uneven surface. (l-m) The same measurements on flat areas of Sb (l) and Cr (m) terminated surfaces. The SFAs are marked by orange arrows. The corresponding sample photos are shown in the bottom left corner of k(i), l(i) and m(i).
Figure 3: Band dispersions of CrSb. (a) The photon energy dependent ARPES spectra along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ measured with photon energies of 97 eV (i), 102 eV (ii), 107 eV (iii), 112 eV (iv), 117 eV (v) and 122 eV (vi) on an uneven surface. (b-c) The same measurements on flat areas of Sb (b) and Cr (c) terminated surfaces. (d-i) The DFT calculated spin polarized band structures along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction at kz = 1 $\pi/c$ (d), 0.8 $\pi/c$ (e), 0.6 $\pi/c$ (f), 0.4 $\pi/c$ (g), 0.2 $\pi/c$ (h) and 0 $\pi/c$ (i) planes. The red bands represent a spin-up and the blue bands spin-down along the $z$ direction. (j-k) Spin-polarized EDCs along the left (j) and right (k) cut lines indicated in [a(ii)]. The spin-polarized EDCs were measured with photon energy of 102 eV, while the red (blue) curve corresponds to spin-up (spin-down) along $z$. The EDCs intensity are already normalized in the non-spin polarized region. In order to obtain sufficient spin statistics, more than 30 hours were accumulated in total. (l) Spin-polarized intensity difference between spin-up and spin-down, with the red (blue) colors indicating spin-up (spin-down) polarization for the left (right) bands.
Figure 4: Weyl fermions and Fermi arcs in CrSb. (a) Calculated total spectral density along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ for CrSb terminated by a Sb (001) surface. (b) surface electronic structure as in (a) but with bulk spectral weight subtracted. The WPs together with the chirality are marked by red ($\chi=1$) and green ($\chi=-1$) dots. Square and round dots represent the opposite-spin WPs at general momenta (WP2) and same-spin WPs (WP3), respectively. (c-g) The photon energy dependent spectra on a Sb terminated surface with photon energies of 77 eV (a), 82 eV (b), 85 eV (c), 89 eV (d), 93 eV (e) under LH polarization along $\rm\overline{M}-\overline{\Gamma}-\overline{M}$. (h-n) The same calculations and measurements as (a-g) for the Cr terminal surface. In (j) on the left also the bulk bands for kz = 0 $\pi/c$ are shown. (o) Calculated total spectral density along $\rm\overline{K}-\overline{\Gamma}-\overline{K}$ of the Sb surface. (p) The same but with bulk contribution subtracted. (q) Measured band dispersion along the $\rm\overline{K}-\overline{\Gamma}-\overline{K}$ direction of a Sb terminated surface with photon energy of 85 eV. (r-t) The same calculations and measurements as (o-q) but on a Cr terminated surface. The bulk calculations for kz = 0.4 $\pi/c$ are also shown on the left side of (q) and (t). The SFAs are marked by orange arrows. The left sides in (c-g) correspond to the EDC second derivative and (j-n) show the MDC second derivative, enhancing the visibility of the bands.